Solve the system.
5x – 3y = 13
4x – 3y = 11

Answer by sdmmadam@yahoo.com(530)

Solve the system.
5x – 3y = 13 ----(1)
4x – 3y = 11 ----(2)
(1)-(2) implies
(5x-4x)=(13-11)
x=2
Put x=2 in (1)
5x-3y = 13
5X2-3y = 13
10-3y=13
10-13=3y
-3=3y
(-3)/3 =y
-1=y
That is y = -1
Answer: x =2 and y = -1
Verification: Putting x =2 and y = -1 in (2)
LHS= 4x-3y = 4X(2)-3X(-1) = 8+3 =11 = RHS
Therefore our values are correct.